# Model-free High Dimensional Mediator Selection with False Discovery Rate Control

### v2

@@ 1 Introduction
To address our data challenges, we develop a novel high-dimensional mediator selection framework with theoretical FDR control guarantees asymptotically. We first frame the mediator selection procedure into a multiple testing problem with composite null hypotheses. Then we use the simultaneous knockoff (Dai & Zheng, 2023) idea to construct the filter statistics for the mediators. The major challenge of this extension is how to handle the unavoidable dependence between path-a and path-b from the exposure variable and potential confounding variables. Our high-level strategy is to decompose path-b to obtain an independent component from path-a.
The main contributions of this paper are listed below:
  1. 1.
    We formulate a general multiple hypothesis testing framework for mediator selection from a high-dimensional candidate set with potentially nonlinear relationships with the exposure or the outcome.
  2. 2.
    For the traditional mediation model, when both path-a and path-b are with linear relationships, we develop a simplified knockoff generating using partial least squares to decompose the path-b to obtain an independent component of path-a.
  3. 3.
    We show the theoretical FDR control guarantees with mild model assumptions for path-b.
  4. 4.
    We perform extensive simulations to demonstrate the FDR control performance and power of our proposed method, and compare with existing methods.
  5. 5.
    We applied our proposed method to analyze the ADNI data.
@@ 2.1 Model settings
We develop our method under a general multiple testing problem for mediator selection (Dai et al., 2024). Let $Y\in\mathbb{R}$ be a one-dimensional response, $X\in\mathbb{R}$ be a one-dimensional exposure , and $\mathbf{M}=(M_{1},\cdots,M_{p})\in\mathbb{R}^{p}$ is the high dimensional mediators, $\mathbf{V}\in\mathbb{R}^{d}$ are low dimensional confounding variables. For $j\in\{1,\cdots,p\}:=[p]$, we define the null hypotheses for path-a and path-b:
@@
$\text{path-b:}~~H_{bj}:Y\perp\!\!\!\perp M_{j}|M_{-j},X,\mathbf{V},$
$\text{where $M_{-j}:=\{M_{k}:k\in[p],k\neq j\}$}.$ (1)
@@
For path-b, (3) can be further relaxed to allow for interactions between the mediators:
@@
where
$\bm{\Delta}=(M_{1}M_{2},\cdots,M_{p-1}M_{p})\in\mathbb{R}^{p(p-1)/2}$
is for two-way interaction terms and $\bm{\delta}$ is its coefficient.
@@ 2.2 Preliminaries
$f($ $[\mathbf{Z}^{1},\widetilde{\mathbf{Z}}^{1}],\ldots,[\mathbf{Z}^{k},\widetilde{\mathbf{Z}}^{k}]_{\text{Swap}(S)},\ldots,[\mathbf{Z}^{K},\widetilde{\mathbf{Z}}^{K}])$
$=f([\mathbf{Z}^{1},\widetilde{\mathbf{Z}}^{1}],\ldots,[\mathbf{Z}^{k},\widetilde{\mathbf{Z}}^{k}],\ldots,[\mathbf{Z}^{K},\widetilde{\mathbf{Z}}^{K}])\odot\epsilon(S),$
@@ 2.3 Mediator Selection algorithm
We design the following procedure to control for the FDR of the selected set of mediators as defined in (5).
@@ 1.2 Path-a test statistics computation
With the above procedures, the final FDR in (5) or a modified FDR
@@ 2.4 Quantification of the mediation effects
We have
NDE $=\mathbb{E}\left[Y(1,\mathbf{M}(0))-Y(0,\mathbf{M}(0))\right],$ (7)
NIE $=\mathbb{E}\left[Y(1,\mathbf{M}(1))-Y(1,\mathbf{M}(0))\right].$ (8)
@@
  • Step 3: Compute $\widehat{\mathbb{E}}\left[Y(x,M_{1}(x_{1}),\ldots,M_{p}(x_{p}))\right]$ by averaging sampled $Y(x,M_{1}(x_{1}),\ldots,M_{p}(x_{p}))$ and average over individual $i$. Then compute $\widehat{\text{NDE}},\widehat{\text{NIE}}$, and $\widehat{\text{NIE}}_{k}$, where
    $\widehat{\text{NDE}}=\widehat{\mathbb{E}}\left[Y(1,\mathbf{M}(0))-Y(0,\mathbf{M}(0))\right],$
    $\widehat{\text{NIE}}=\widehat{\mathbb{E}}\left[Y(1,\mathbf{M}(1))-Y(1,\mathbf{M}(0))\right],\text{ and}$
@@ 2.5 Theoretical results
The proof of Theorem 1 is postponed to Web Appendix B.
In Theorem 1 we provide the theoretical guarantee for the algorithm in Section 2.4 under the assumption of the large number of weakly dependent path-a statistics.
@@ 3.1 Data Generation and Simulation Settings
For each subject $i\in[n]$, the exposure variable $X_{i}\in\mathbb{R}$ and the covariate $V_{i}\in\mathbb{R}$ are generated from a multivariate normal distribution, with mean $\begin{pmatrix}0\\ 0\end{pmatrix}$ and variance-covariance matrix $\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix}$. The correlated potential mediators $\mathbf{M}_{i}=(M_{i1},\dots,M_{ip})^{\top}$ is generated using $\mathbf{M}_{i}=X_{i}\bm{\alpha}+V_{i}\bm{\eta}_{1}+\bm{\epsilon}_{1i}$, where $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{p})^{\top}$$\bm{\eta_{1}}=(\eta_{11},\dots,\eta_{1p})^{\top}$ and $\bm{\epsilon}_{1i}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma})$, where $\bm{\Sigma}$ is a variance-covariance matrix with compound symmetry structure (i.e.,diagonal entries are 1 and off-diagonal entries are $\rho$). The outcome $Y_{i}$ is generated as $Y_{i}=X_{i}\gamma+\mathbf{M}^{\top}\bm{\beta}+V_{i}\eta_{2}+\epsilon_{2i}$, where $\bm{\epsilon}_{2i}\sim\mathcal{N}(0,\sigma^{2})$. Detailed simulation settings for parameters $\bm{\alpha},\bm{\eta}_{1},\rho,\gamma,\bm{\beta},\eta_{2},\sigma$ are provided in Appendix C.1.
@@
We consider an scenario where path-a remain the same as the linear model setting; while in path-b, only interaction effects (i.e, interaction between mediators) are present. The generation of the exposure variable $X_{i}$ and potential mediators $\mathbf{M}_{i}$ are the same as the linear setting. The outcome $Y_{i}$ is generated as $Y_{i}=X_{i}\gamma+Z_{i}\bm{\delta}+\epsilon_{2i}$, where $Z_{i}=(M_{i1}M_{i2},\dots,M_{i(p-1)}M_{ip})$, $\epsilon_{2i}\overset{i.i.d}{\sim}\mathcal{N}(0,\sigma^{2})$. Detailed simulation settings for parameters $\bm{\alpha},\bm{\eta}_{1},\rho,\gamma,\bm{\delta},\sigma^{2}$ are provided in Appendix C.1.
@@
For this setting, the generation of the exposure variable $X_{i}$ and potential mediators $\mathbf{M}_{i}$ are still the same as the linear model setting. We extend our path-b model to a nonlinear model by applying the cosine function to each element of $\mathbf{M}$. Specifically, in path-b, the outcome $Y$ is generated as $Y_{i}=X_{i}\gamma+\sum_{j=1}^{p}\beta_{j}\cdot cos(M_{ij})+\epsilon_{2i}$, where $\epsilon_{2i}\overset{i.i.d}{\sim}\mathcal{N}(0,\sigma^{2})$. Detailed simulation settings for parameters $\bm{\alpha},\bm{\eta}_{1},\rho,\gamma,\bm{\beta},\sigma^{2}$ are provided in Appendix C.1.
Further details on the data generation process are also provided in Appendix C.1.
3.2 Method specification
[…]
@@ 4 Data Analysis
To obtain the structural MRI data, we first convert ADNI dataset to follow the Brain Imaging Data Structure (BIDS) rule via Clinica (Routier et al., 2021) and applied fMRIPrep (Esteban et al., 2019) with the FreeSurfer (fischl2012freesurfer) reconstruction workflow to pre-process the data. This pipeline outputs 555 structural MRI-derived measures of each subjects (e.g. cortical thickness, surface area and brain volumes) corresponding to specific brain regions or morphometric features, such as the average cortical thickness in various gyri and sulci, and volumetric data for subcortical structures. The full list of the 555 structural MRI-derived measures can be found in Web Appendix D.1.
@@
The outcome of interest in this study is the change in multiple cognitive assessment scores from baseline to 17 years of follow-up. These assessments are typically conducted at baseline and repeated at regular intervals (e.g., every 3 or 6 months). To account for the missing visits, we apply the last observation carried forward (LOCF) method (Lachin, 2016) to impute missing cognitive scores. To evaluate different aspects of cognitive function, we apply multiple cognitive scores, including the Clinical Dementia Rating Sum of Boxes (CDR-SB), Alzheimer’s Disease Assessment Scale-Cognitive Subscale (ADAS-Cog) and Mini-Mental State Examination (MMSE). Higher CDR-SB and ADAS-Cog scores, and lower MMSE scores, are each indicative of greater Alzheimer’s disease severity. More details about those cognitive scores can be found in Appendix D.2.
@@
Finally, we apply our proposed GKMS method to identify important structural MRI-derived measures from a set of $p=555$ candidate correlated mediators that may mediate the relationship between gender and changes in cognitive scores among $n=1097$ MCI patients. We consider Age at baseline as a confounding variable $V$. The FDR is set as $q=0.2$. For exposure X, we set males as the reference group. $\mathbf{Y}$ are the changes from baselines in the cognitive assessment scores CDR-SB, ADAS-Cog and MMSE. Both the outcome $\mathbf{Y}$ and the mediators $\mathbf{M}$ are scaled without centering ensuring that the values the $\alpha$ and $\beta$ are the standardized regression coefficients. Tables 3, 4 and 5 summarized all the mediator selection results with their estimated mediation effect for changes in CDR-SB, ADAS-Cog and MMSE scores, respectively. For our proposed method with random forest estimation in path-b (GKMS (RF)), we evaluate the mediator effect using g-formula with numerical Monte Carlo integration as described by Imai et al. (2010), For the other methods, we employed structural equation modeling (SEM) through the lavaan R package (Rosseel, 2012). More details on the estimation of mediation effects are provided in Appendix D.3.
@@
These findings are consistent with previous independent studies. For example, A longitudinal analysis found that women with probable Alzheimer’s disease exhibited significantly faster hippocampal atrophy compared to men (Ardekani et al., 2016). This accelerated atrophy was associated with a more rapid decline into dementia, highlighting female sex as a risk factor for faster disease progression. Another study analyzing resting-state fMRI data from individuals with mild cognitive impairment (MCI) showed that females had weaker hippocampal connectivity to the precuneus cortex and brainstem compared to males. This diminished connectivity may contribute to the increased susceptibility of women to cognitive decline and progression to Alzheimer’s disease (Williamson et al., 2022). Poulin et al. (2011) reported that amygdala atrophy is associated with increased dementia severity. Males consistently exhibit larger amygdala volumes than females, after adjusting for total brain volume (Goldstein et al., 2001). Moreover, Tau-PET studies have shown that tau accumulation in the amygdala surpasses that in the entorhinal cortex, indicating the amygdala may be among the first brain regions affected by tau pathology in Alzheimer’s disease (Stouffer et al., 2024; Lupton et al., 2016). Cortical thinning in the Angular Gyrus and superior temporal sulcus have also been frequently linked to AD and other neurodegenerative disorders (Dickerson et al., 2009). Our analysis provides insights into how gender influences cognitive decline and contributes to a better understanding of gender-specific differences in dementia progression.
@@ 5 Discussions
Method Selected Mediator Effect (95% CI; p-value)
GKMS (RF)* Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; <<0.001)
@@
Thick Avg of Angular Gyrus (L) $\alpha$ = 0.23 (0.115, 0.346; <<0.001)
@@
$\beta$ = -0.132 (-0.193; -0.071;<<0.001)
Mediation = 0.069 (0.034, 0.104; <<0.001)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.178 (-0.238; -0.118;<<0.001)
Mediation = 0.09 (0.054, 0.126; <<0.001)
Volume of Amygdala (R) $\alpha$ = -0.596 (-0.706, -0.487; <<0.001)
$\beta$ = -0.111 (-0.171; -0.052;<<0.001)
Mediation = 0.066 (0.029, 0.104; 0.001)
$\beta$ = 0.15 (0.089; 0.211;<<0.001)
Mediation = -0.072 (-0.106, -0.039; <<0.001)
Volume of Brain Stem $\alpha$ = -0.955(-1.061, -0.848; <<0.001)
$\beta$ = -0.121 (0.06; 0.183;<<0.001)
Mediation = -0.116 (-0.176, -0.056; <<0.001)
$\beta$ = -0.376 (-0.438; -0.314;<<0.001)
Mediation = 0.19 (0.139, 0.241; <<0.001)
HIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.282 (-0.343, -0.222; <<0.001)
Mediation = 0.143 (0.099, 0.186; <<0.001)
Volume of Inf.Lat.Vent (R) $\alpha$ = -0.483(-0.591, -0.376; <<0.001)
$\beta$ = 0.135 (0.074, 0.196; <<0.001)
Mediation = -0.065 (-0.098, -0.032; <<0.001)
ThickAvg of Parahip2 (R) $\alpha$ = 0.123 (0.007, 0.24; 0.037)
$\beta$ = -0.128 (-0.185, -0.072; <<0.001)
Mediation = -0.016 (-0.032, 0.001; 0.059)
$\beta$ = -0.376 (-0.438; -0.314;<<0.001)
Mediation = 0.19 (0.139, 0.241; <<0.001)
  • *For GKMS (RF) method, $\beta$ effects are the importance scores with direction derived from partial dependence analysis comparing outcome predictions at the first and third quartiles of mediator values. Mediation effects are estimated using Monte Carlo integration as described by Imai et al. (2010) and confidence intervals are obtained via bootstrap.
  • 1 Inf.Lat.Vent refers to the Inferior Lateral Ventricle.
  • 2 Parahip represents the medial temporal-occipital parahippocampal gyrus.
Method Selected Mediator Effect (95% CI; p-value)
GKMS (RF)* Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; <<0.001)
@@
ThickAvg of S.temporal.sup1(R) $\alpha$ = 0.092 (-0.023, 0.207; 0.117)
@@
Volume of white matter in ITL 2(L) $\alpha$ = -0.895 (-1.004, -0.785; <<0.001)
@@
$\beta$ = -0.083 (-0.145, -0.02; 0.009)
Mediation = 0.043 (0.009, 0.077; 0.012)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.139 (-0.201, -0.078; <<0.001)
Mediation = 0.07 (0.036, 0.105; <<0.001)
ThickAvg of Angular Gyrus(L) $\alpha$ = 0.233 (0.118, 0.349; <<0.001)
$\beta$ = -0.099 (-0.157, -0.041; 0.001)
Mediation = -0.023 (-0.041, -0.005; 0.011)
ThickAvg of S.temporal.sup(R) $\alpha$ = 0.092 (-0.023, 0.207; 0.117)
$\beta$ = -0.167 (-0.226, -0.109; <<0.001)
Mediation = -0.015 (-0.035, 0.005; 0.131)
$\beta$ = 0.081 (0.019, 0.143; 0.011)
Mediation = -0.058 (-0.104, -0.013; 0.012)
HIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.324 (-0.387, -0.261; <<0.001)
Mediation = 0.164 (0.116, 0.211; <<0.001)
$\beta$ = -0.324 (-0.387, -0.261; <<0.001)
Mediation = 0.164 (0.116, 0.211; <<0.001)
$\beta$ = -0.063 (-0.123, -0.004; 0.036)
Mediation = 0.013 (-0.001, 0.027; 0.078)
ThickAvg of Pole temporal(R) $\alpha$ = 0.259 (0.143, 0.376; <<0.001)
$\beta$ = -0.084 (-0.146, -0.023; 0.007)
Mediation = -0.022 (-0.041, -0.003; 0.022)
  • We use version ADAS-13 representing the results of ADAS-Cog.
  • *For GKMS (RF) method, $\beta$ effects are the importance scores with direction derived from partial dependence analysis comparing outcome predictions at the first and third quartiles of mediator values. Mediation effects are estimated using Monte Carlo integration as described by Imai et al. (2010) and confidence intervals are obtained via bootstrap.
  • 1 S.temporal.sup refers to the average cortical thickness of the region superior temporal sulcus from the aparc.a2009s statistics file.
  • 2 ITL means inferior temporal lobe. This measure represents the white matter in the left hemisphere of the inferior temporal lobe region.
Method Selected Mediator Effect (95% CI; p-value)
GKMS (RF)* Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; <<0.001)
@@
ThickAvg of G-front-middle1 (L) $\alpha$ = 0.29 (0.175, 0.404; <<0.001)
@@
$\beta$ = 0.111 (0.048, 0.174; 0.001)
Mediation = -0.058 (-0.093, -0.023; 0.001)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = 0.184 (0.122, 0.246; <<0.001)
Mediation = -0.093 (-0.13, -0.056; <<0.001)
Volume of Amygdala (R) $\alpha$ = -0.596 (-0.706, -0.487; <<0.001)
$\beta$ = 0.068 (0.007, 0.129; 0.028)
Mediation = -0.041 (-0.078, -0.004; 0.032)
Volume of Inf.Lat.Vent2 (R) $\alpha$ = -0.483(-0.591, -0.376; <<0.001)
$\beta$ = -0.088 (-0.15, -0.026; 0.006)
Mediation = 0.042 (0.011, 0.074; 0.008)
$\beta$ = -0.147 (-0.206, -0.088; <<0.001)
Mediation = 0.103 (0.059, 0.148; <<0.001)
$\beta$ = 0.336 (0.273, 0.399; <<0.001)
Mediation = -0.17 (-0.218, -0.121; <<0.001)
HIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = 0.219 (0.158, 0.281; <<0.001)
Mediation = -0.111 (-0.15, -0.072; <<0.001)
ThickAvg of S.temporal.sup3 (L) $\alpha$ = 0.12 (0.007, 0.233; 0.037)
$\beta$ = 0.246 (0.187, 0.305; 0)
Mediation = 0.03 (0.001, 0.058; 0.043)
$\beta$ = 0.336 (0.273, 0.399; <<0.001)
Mediation = -0.17 (-0.218, -0.121; <<0.001)
  • *For GKMS (RF) method, $\beta$ effects are the importance scores with direction derived from partial dependence analysis comparing outcome predictions at the first and third quartiles of mediator values. Mediation effects are estimated using Monte Carlo integration as described by Imai et al. (2010) and confidence intervals are obtained via bootstrap.
  • 1 G-front-middle refers the middle frontal gyrus, an anatomical brain region.
  • 2 Inf.Lat.Vent refers to the Inferior Lateral Ventricle.
  • 3 ThickAvg means the average cortical thickness. S.temporal.sup refers to superior temporal sulcus. This measure means the average cortical thickness of the region superior temporal sulcus from the aparc.a2009s statistics file.
@@
Lemma B3.
Assume $p_{j}$s satisfy the requirement in Lemma B2, then with the Knockoff filter, the mFDR can be controlled at level $q$ and with the Knockoff+ filter, the FDR can be controlled at level $q$ when $p\rightarrow\infty$.
@@
where the first inequality holds by the definition of $\widehat{k}_{+}$ and the second inequality holds by Lemma B2.
@@
where the first inequality holds by the definition of $\widehat{k}_{0}$ and the second inequality holds by Lemma B2. ∎

### v3

@@ 1 Introduction
To address our data challenges, we develop a novel high-dimensional mediator selection framework with theoretical FDR control guarantees asymptotically. We first frame the mediator selection procedure into a multiple testing problem with composite null hypotheses. Then we use the simultaneous knockoff (Dai & Zheng, 2023) idea to construct the filter statistics for the mediators. The major challenge of this extension is how to handle the unavoidable dependence between path-a and path-b from the exposure variable and potential confounding variables. Our high-level strategy is to decompose path-b to obtain an independent component from path-a. The main contributions of this paper are listed below:
1. We formulate a general multiple hypothesis testing framework for mediator selection from a high-dimensional candidate set with potentially nonlinear relationships with the exposure or the outcome.
2. For the traditional mediation model, when both path-a and path-b are with linear relationships, we develop a simplified knockoff generating using partial least squares to decompose the path-b to obtain an independent component of path-a.
3. We show the theoretical FDR control guarantees with mild model assumptions for path-b.
4. We perform extensive simulations to demonstrate the FDR control performance and power of our proposed method, and compare with existing methods.
5. We applied our proposed method to analyze the ADNI data.
@@ 2.1 Model settings
We develop our method under a general multiple testing problem for mediator selection (Dai et al., 2024). Let $Y\in\mathbb{R}$ be a one-dimensional response, $X\in\mathbb{R}$ be a one-dimensional exposure , and $\mathbf{M}=(M_{1},\cdots,M_{p})\in\mathbb{R}^{p}$ is the high dimensional mediators, $\mathbf{V}\in\mathbb{R}^{d}$ are low dimensional confounding variables. For $j\in\{1,\cdots,p\}:=[p]$, we define the null hypotheses for path-a: $H_{aj}$ and path-b: $H_{bj}$:
@@
$\text{path-b:}~~H_{bj}:Y\perp\!\!\!\perp M_{j}|M_{-j},X,\mathbf{V},\text{where $M_{-j}:=\{M_{k}:k\in[p],k\neq j\}$}.$
@@
For path-b, equation (2) can be relaxed to allow for interactions between the mediators:
@@
where $\bm{\Delta}=(M_{1}M_{2},\cdots,M_{p-1}M_{p})\in\mathbb{R}^{p(p-1)/2}$ is for two-way interaction terms and $\bm{\delta}$ is its coefficient.
@@ 2.2 Preliminaries
$f($ $[\mathbf{Z}^{1},\widetilde{\mathbf{Z}}^{1}],\ldots,[\mathbf{Z}^{k},\widetilde{\mathbf{Z}}^{k}]_{\text{Swap}(S)},\ldots,[\mathbf{Z}^{K},\widetilde{\mathbf{Z}}^{K}])=f([\mathbf{Z}^{1},\widetilde{\mathbf{Z}}^{1}],\ldots,[\mathbf{Z}^{k},\widetilde{\mathbf{Z}}^{k}],\ldots,[\mathbf{Z}^{K},\widetilde{\mathbf{Z}}^{K}])\odot\epsilon(S),$
@@ 2.3 Mediator Selection algorithm
We design the following procedure to control for the FDR of the selected set of mediators as defined in equation (4).
@@ 1.2 Path-a test statistics computation
With the above procedures, the final FDR in equation (4) or a modified FDR
@@ 2.4 Quantification of the mediation effects
We have $NDE=\mathbb{E}Y(1,\mathbf{M}(0))-Y(0,\mathbf{M}(0))$, $NIE=\mathbb{E}Y(1,\mathbf{M}(1))-Y(1,\mathbf{M}(0))$.
@@
Step 3: Compute $\widehat{\mathbb{E}}\left[Y(x,M_{1}(x_{1}),\ldots,M_{p}(x_{p}))\right]$ by averaging sampled
$Y(x,M_{1}(x_{1}),\ldots,M_{p}(x_{p}))$ and average over individual $i$. Then compute $\widehat{\text{NDE}},\widehat{\text{NIE}}$, and $\widehat{\text{NIE}}_{k}$:
$\widehat{\text{NDE}}=\widehat{\mathbb{E}}\left[Y(1,\mathbf{M}(0))-Y(0,\mathbf{M}(0))\right],\widehat{\text{NIE}}=\widehat{\mathbb{E}}\left[Y(1,\mathbf{M}(1))-Y(1,\mathbf{M}(0))\right],\text{ and}$
@@ 2.5 Theoretical results
The proof of Theorem 1 is postponed to Web Appendix B. In Theorem 1 we provide the theoretical guarantee for the algorithm in Section 2.3 under the assumption of the large number of weakly dependent path-a statistics.
@@ 3.1 Data Generation and Simulation Settings
For each subject $i\in[n]$, the exposure variable $X_{i}\in\mathbb{R}$ and the covariate $V_{i}\in\mathbb{R}$ are generated from a multivariate normal distribution, with mean $\begin{pmatrix}0\\ 0\end{pmatrix}$ and variance-covariance matrix $\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix}$. The correlated potential mediators $\mathbf{M}_{i}=(M_{i1},\dots,M_{ip})^{\top}$ is generated using $\mathbf{M}_{i}=X_{i}\bm{\alpha}+V_{i}\bm{\eta}_{1}+\bm{\epsilon}_{1i}$, where $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{p})^{\top}$$\bm{\eta_{1}}=(\eta_{11},\dots,\eta_{1p})^{\top}$ and $\bm{\epsilon}_{1i}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma})$, where $\bm{\Sigma}$ is a variance-covariance matrix with compound symmetry structure (i.e.,diagonal entries are 1 and off-diagonal entries are $\rho$). The outcome $Y_{i}$ is generated as $Y_{i}=X_{i}\gamma+\mathbf{M}^{\top}\bm{\beta}+V_{i}\eta_{2}+\epsilon_{2i}$, where $\bm{\epsilon}_{2i}\sim\mathcal{N}(0,\sigma^{2})$.
@@
We consider an scenario where path-a remain the same as the linear model setting; while in path-b, only interaction effects (i.e, interaction between mediators) are present. The generation of the exposure variable $X_{i}$ and potential mediators $\mathbf{M}_{i}$ are the same as the linear setting. The outcome $Y_{i}$ is generated as $Y_{i}=X_{i}\gamma+Z_{i}\bm{\delta}+\epsilon_{2i}$, where $Z_{i}=(M_{i1}M_{i2},\dots,M_{i(p-1)}M_{ip})$, $\epsilon_{2i}\overset{i.i.d}{\sim}\mathcal{N}(0,\sigma^{2})$.
@@
For this setting, the generation of the exposure variable $X_{i}$ and potential mediators $\mathbf{M}_{i}$ are still the same as the linear model setting. We extend our path-b model to a nonlinear model by applying the cosine function to each element of $\mathbf{M}$. Specifically, in path-b, the outcome $Y$ is generated as $Y_{i}=X_{i}\gamma+\sum_{j=1}^{p}\beta_{j}\cdot cos(M_{ij})+\epsilon_{2i}$, where $\epsilon_{2i}\overset{i.i.d}{\sim}\mathcal{N}(0,\sigma^{2})$.
Details of the data generation process and parameter settings are provided in Appendix C.1, and method specifications in Appendix C.2.
@@ 4 Data Analysis
To obtain the structural MRI data, we first convert ADNI dataset to follow the Brain Imaging Data Structure (BIDS) rule via Clinica (Routier et al., 2021) and applied fMRIPrep (Esteban et al., 2019) with the FreeSurfer (Fischl, 2012) reconstruction workflow to pre-process the data. This pipeline outputs 555 structural MRI-derived measures of each subjects (e.g. cortical thickness, surface area and brain volumes) corresponding to specific brain regions or morphometric features, such as the average cortical thickness in various gyri and sulci, and volumetric data for subcortical structures. The full list of the 555 structural MRI-derived measures can be found in Web Appendix D.1.
@@
The outcome of interest in this study is the change in multiple cognitive assessment scores from baseline to 17 years of follow-up. These assessments are typically conducted at baseline and repeated at regular intervals (e.g., every 3 or 6 months). To account for the missing visits, we apply the last observation carried forward (LOCF) method (Lachin, 2016) to impute missing cognitive scores. To evaluate different aspects of cognitive function, we apply multiple cognitive scores, including the Clinical Dementia Rating Sum of Boxes (CDR-SB), Alzheimer’s Disease Assessment Scale-Cognitive Subscale (ADAS-Cog) and Mini-Mental State Examination (MMSE). Higher CDR-SB and ADAS-Cog scores, and lower MMSE scores, each indicate greater AD severity. More details about those cognitive scores can be found in Appendix D.2.
@@
We apply the proposed GKMS method to identify important structural MRI-derived measures from a set of $p=555$ candidate correlated mediators that may mediate the relationship between gender and changes in cognitive scores among $n=1097$ MCI patients. We consider Age at baseline as a confounding variable $V$. The FDR is set as $q=0.2$. For exposure X, we set males as the reference group. $\mathbf{Y}$ are the changes from baselines in the cognitive assessment scores CDR-SB, ADAS-Cog and MMSE. Both the outcome $\mathbf{Y}$ and the mediators $\mathbf{M}$ are scaled, ensuring that the values the $\alpha$ and $\beta$ are the standardized regression coefficients. To capture the interactions among the mediators, we applied our proposed GKMS method with Random Forests in path-b. Table 3 summarizes the mediator selection results with their estimated mediation effects for changes in CDR-SB, ADAS-Cog, and MMSE scores. The mediator effects were estimated using the g-formula with numerical Monte Carlo integration as described by Imai et al. (2010). More details on the estimation of mediation effects are provided in Appendix D.3.
Our proposed method identified both shared and distinct mediators across the three cognitive outcome measures. The hippocampal and amygdala volumes were consistently selected as important mediators for all three cognitive outcomes. Cortical thickness in the left angular gyrus was chosen as a significant mediator for both CDR-SB and ADAS-Cog. Furthermore, the cortical thickness in the middle frontal gyrus (left) was also selected as a positive mediator for MMSE.
As a sensitivity analysis, we applied our proposed method with Lasso linear regression in path-b (GKMS (Lasso)), HIMA, HIMA2, eHIMA, and MediationFDR methods for comparison. The mediation effects were estimated using structural equation modeling (SEM) through the lavaan R package (Rosseel, 2012). Tables 4, 5, and 6 summarize the results for CDR-SB, ADAS-Cog, and MMSE scores, respectively. Both hippocampal and amygdala volumes are commonly selected in those methods. More details are provided in Appendix D.4.
@@
These findings are consistent with previous independent studies. For example, a longitudinal analysis found that women with probable Alzheimer’s disease exhibited significantly faster hippocampal atrophy compared to men (Ardekani et al., 2016). This accelerated atrophy was associated with a more rapid decline into dementia, highlighting female sex as a risk factor for faster disease progression. Another study analyzing resting-state fMRI data from individuals with mild cognitive impairment (MCI) showed that females had weaker hippocampal connectivity to the precuneus cortex and brainstem compared to males. This diminished connectivity may contribute to the increased susceptibility of women to cognitive decline and progression to Alzheimer’s disease (Williamson et al., 2022). Poulin et al. (2011) reported that amygdala atrophy is associated with increased dementia severity. Males consistently exhibit larger amygdala volumes than females, after adjusting for total brain volume (Goldstein et al., 2001). Moreover, Tau-PET studies have shown that tau accumulation in the amygdala surpasses that in the entorhinal cortex, indicating the amygdala may be among the first brain regions affected by tau pathology in Alzheimer’s disease (Stouffer et al., 2024; Lupton et al., 2016). Our analysis provides insights into how gender influences cognitive decline and contributes to a better understanding of gender-specific differences in dementia progression.
@@ 5 Discussions
Table 3: ADNI study mediator selection results using GKMS method for the three outcomes: CDR-SB, ADAS-Cog, and MMSE.(L): Left; (R): Right.
Cognitive Scores Selected Mediator Effect (95% CI; p-value)*
CDR-SB Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; <<0.001)
@@
ThickAvg1 of Angular Gyrus (L) $\alpha$ = 0.23 (0.115, 0.346; <<0.001)
@@
ADAS-Cog Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; <<0.001)
@@
ThickAvg of S.temporal.sup2 (R) $\alpha$ = 0.092 (-0.023, 0.207; 0.117)
@@
Volume of white matter in ITL3 (L) $\alpha$ = -0.895 (-1.004, -0.785; <<0.001)
@@
MMSE Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; <<0.001)
@@
ThickAvg of G-front-middle4 (L) $\alpha$ = 0.29 (0.175, 0.404; <<0.001)
@@
  • $\beta$ effects are the importance scores with direction derived from partial dependence analysis comparing outcome predictions at the first and third quartiles of mediator values. Mediation effects are estimated using Monte Carlo integration as described by Imai et al. (2010), and confidence intervals are obtained via bootstrap.
  • Abbreviations (all extracted from the aparc.a2009s statistics file):
  • 1 ThickAvg: Average cortical thickness of the specified cortical region.
  • 2 S.temporal.sup: Superior temporal sulcus region.
  • 3 ITL: Inferior temporal lobe, denoting white matter volume in the left hemisphere of the inferior temporal lobe.
  • 4 G-front-middle: Middle frontal gyrus, a cortical brain region.
@@
Lemma 3.
Assume $p_{j}$s satisfy the requirement in Lemma 2, then with the Knockoff filter, the mFDR can be controlled at level $q$ and with the Knockoff+ filter, the FDR can be controlled at level $q$ when $p\rightarrow\infty$.
@@
where the first inequality holds by the definition of $\widehat{k}_{+}$ and the second inequality holds by Lemma 2.
@@
where the first inequality holds by the definition of $\widehat{k}_{0}$ and the second inequality holds by Lemma 2. ∎